System

Date Time Event Comments / Remarks
RAL 19 JULY 2007

17:45

RAL-LCG2 ralsrmb.rl.ac.uk SE down 20th July 2007 planned to shut down access to the Atlas Castor instance tomorrow 20th July at 8.30AM UTC
CERN 19 JULY 2007

13:57

CERN Prod FTS - Intervention, 08:00, July 20th Duration: One hour Impact: Service Dedgregation
IN2P3 19 JULY 2007

13:09

Scheduled downtime at IN2P3-LPC and AUVERGRID Scheduled downtime planed from Monday, 23 July at 4H00 AM to Tuesday, 24 July at 12H00 AM.
SARA 17 JULY 2007

09:12

Unscheduled intervention at SARA Problems with pnfs and set the CERN-SARA FTS channel to be Inactive.
NDGF 16 JULY 2007

15:44

NDGF downtime on thursday 19th July srm.ndgf.org will be down from 07:00 UTC  to 18:00 UTC
INFN 16 JULY

13:33

Roma1 storage temporary unavailable DPM storage will be unavailable untill 18-july 17.00 UTC
FZK/GridKa  14 JULY 2007   FZK/GridKa Mass storage (tape) connectivity restored  
BNL 13 JULY 2007   BNL 3D Oracle Cluster Transparent intervention 18 Jul 2007 between 13:30-16:30 UTC time  
RAL 12 JULY 2007

13:13

Minor intervention to GOCDB3 web interface The intervention will take place at 12.30PM BST (11.30AM UTC) Friday July 13th 2007.
CNAF 11 JULY 2007

17:20

CASTOR upgrade at CNAF is over CASTOR@CNAF has been upgraded to version 2.1.3-15
CERN 11 JULY 2007

09:49  

CERN prod-bdii problem solved It has been upgraded
CERN 10 JULY 2007

13:37 

Downgrade of prod-bdii at CERN A problem under investigation has caused a downtime of CERN-PROD
CERN 09 JULY 2007

16:06

Scheduled network intervention at CERN on Wed, July 18th 7.00am UTC The CERN services which may be impacted are:  FTS ,LFC,  VOMS/VOMRS (ALICE, ATLAS, CMS, LHCb, DTEAM, OPS, Sixt, Unosat, Geant4),  SAM and GridView
CERN 08 JULY 2007

16:56

Status of Castor - Sun 14:30 Castorpublic services are now restored, and data transfers have resumed.
CNAF 06 JULY 2007   INFN-CNAF RBs and WMS closed (submissions forbidden) for glite updates 26/27/28 From this afternoon 17:00 UTC until Monday

 

CNAF 05 JULY 2007

15:17

Service intervention at CNAF(09.07.2007 10:00 - 09.07.2007 14:00) LHCb-LFC database will be affected
RAL 05 JULY 2007

12:49

GOCDB3 Urgent database maintenance the NGS Oracle cluster hosting GOCDB3 will be rebooted tonight (5th July 2007) at 23:00 BST
CERN 05 JULY 2007

11:52 

CERN-PROD disappeared from BDII => SAM fake alarms Due to a site-BDII problem
NDGF 05 JULY 2007

09:47

NDGF-T1 SRM available  
NDGF 05 JULY 2007

09:08

NDGF-T1 SRM service unavailable Troubleshooting in progress. Estimated time of recovery by 10.00 UTC
CERN 04 JULY 2007

15:01

CERN myproxy FTS problem -- solved  
RAL 04 JULY 2007

13:18

Accounting - Unscheduled Downtime The accounting server at RAL is down in order to address a filesystem issue
CERN 04 JULY 2007

13:18

CERN myproxy FTS problem CERN FTS unavailable since this morning. Reason: the host certificate of the myproxy server for FTS transfers expired yesterday
CERN 04 JULY 2007

13:02

CERN CASTOR problem Since last night, suffering from intermitent problems
CNAF 04 JULY 2007

12:24

CASTOR downtime at CNAF from the 9th to the 11th of July The intervention will start at 9:00 UTC of monday 9 july and will end at 18:00 UTC of wednesday 11 july.
RAL 03 JULY 2007 

11:26

RAL-LCG2 RB downtime finished The LCG Resource Broker lcgrb01.gridpp.rl.ac.uk is now back into production
RAL 02 JULY 2007

16:53

GOCDB urgent intervention this evening NGS Oracle database at RAL, which hosts GOCDB3, requires urgent mainentance. Urgent work will be done tonight, 02/07/2007, at 23:00 BST
RAL 02 JULY 2007

13:54

Minor intervention to GOCDB3 web interface Intervention will take place at 10AM UTC Tuesday July 3rd 2007.

 

SARA 02 JULY 2007

09:59

Service intervention at SARA(02.07.2007 07:30 - 02.07.2007 08:30) saradb will be affected
GRIDKA 29 JUNE 2007   FZK/GridKa downtime ended  
SARA 29 JUNE 2007  

16:01

Service intervention at SARA (29.06.2007 14:00 - 29.06.2007 14:30) saradb will be affected  
SARA 29 JUNE 2007  

13:17

Service intervention at SARA(29.06.2007 09:30 - 29.06.2007 11:00) saradb will be affected 
GRIDKA 28 JUNE 2007

21:43

Extended scheduled down time for FZK/GridKa till 29/6 10:00 UTC  
GRIDKA 28 JUNE 2007

13:05

GGUS system will be down from 13:00 to 16:00 UTC. New release of the GGUS portal will be made available
IN2P3 28 JUNE 2007

12:28

CIC portal successfully released CIC portal is back online. Current version is 4.4

 

SARA 28 JUNE 2007  

11:11

Service intervention at SARA (27.06.2007 22:00 - 28.06.2007 09:00) saradb will be affected
PIC 27 JUNE 2007

14:14

FTS service at PIC is back again  
SARA 27 JUNE 2007 

10:40

MSS problem at SARA is again  
TRIUMP 27 JUNE 2007

02:56

Network issues at TRIUMF 1 Gb failover lightpath is down because of recurring problems with 2 links.  5 Gb lightpath down as part of a previously scheduled maintenance in eastern Canada (a 3 hour window)
SARA 26 JUNE 2007

11:41

MSS problem at SARA is solved  
SARA 26 JUNE 2007

09:37

Problems with MSS at SARA  
PIC 25 JUNE 2007

16:30

FTS service at pic is down Due to a problem with the fts software
CERN 25 JUNE 2007

 13:52

CASTORALICE upgrade on Wednesday, June 27. ALICE Castor instance at CERN will be upgraded.  Intervention to take a whole working day starting at 9:00 a.m. (Geneva time)  
RAL 22 JUNE 2007   RAL-LCG2 OPN link restored  
CERN 22 JUNE 2007   CDB intervention Monday June 25th 17h30 - has been deferred to Thursday, June 28th at 17h30  
IN2P3 22 JUNE 2007

14:32

CIC portal release and downtime on June 28th (unavailable from 7:00 UTC   to 10:00 UTC ) Next version of the CIC portal will be released on Thursday, June 28th.  
GRIDKA 21 JUNE 2007

17:54

Service intervention at GRIDKA (27.06.2007 05:00 - 28.06.2007 18:00 UTC) All DBs will be affected
CERN 21 JUNE 2007

12:32

SAM web services were not available since 20th June evening ~20:00 until 21th June morning ~10:00.  
SARA 21 JUNE 2007

11:40

Service intervention at SARA (20.06.2007 07:40 - 21.06.2007 08:00) saradb will be affected
CNAF 21 JUNE 2007

11:30

Service intervention at CNAF(21.06.2007 15:00 - 21.06.2007 18:00 UTC) Due to an high priority database maintenance
CERN 21 JUNE 2007

09:06 

SE failure on castorgridsc.cern.ch -(CERN-PROD)-  
CERN 21 JUNE 2007

05:55 

SRM failure on srm.cern.ch -(CERN-PROD)-  
SARA 20 JUNE 2007

15:11

dCache upgrade @ SARA is complete  
PIC 20 JUNE 2007

14:52

srm-disk service downtime at pic 21-22 June Due to an upgrade of the underlying dcache infrastructure to 1.7
CERN 20 JUNE 2007

13:53 

The CERN T0 export FTS prod-fts-ws.cern.ch is  back in production  
SARA 20 JUNE 2007

13:03

dCache upgrade @ SARA  
RAL 20 JUNE 2007

12:41

OPN link between RAL-LCG2 and CERN down  
CERN 19 JUNE 2007

03:26

CERN-PROD FTS prod-fts-ws.cern.ch down Wednesday 20th Morning  
RAL 19 JUNE 2007

12:46

RAL-LCG2 - lcgrb01 unscheduled maintenance  
CERN 18 JUNE 2007

22:31

Downtime of ce109 and ce112 to fix a problem Downtime is expected to end on 19.06.2007 afternoon at 17h00

 

CERN 18 JUNE 2007

21:54

FTS failure on prod-fts-ws.cern.ch -(CERN-PROD)-  
ASGC 18 JUNE 2007

21:02

ASGC cms queue stop during CASTOR-2 intervention on 20.06.2007 ASGC is going to re-organize disk servers and upgrade CASTOR-2 to newer version  
RAL 18 JUNE 2007 

18:17

RAL-LCG2 Tape services unavailable 26th June 7-11am GMT To allow network switches to be moved into a dedicated rack
GRIDKA 18 JUNE 2007 

16:03

Scheduled maintenance at FZK/GridKa from 27/6 05:00 - 28/6 18:00(UTC)  
CERN/BNL 18 JUNE 2007

11:15

Traffic interruption between BNL and CERN on Jun-19-2007 (17:00 to 18:00 CET) Areas Affected: BNL / CERN Internet traffic

 

TRIUMP 15 JUNE 2007   Service intervention at TRIUMF  (18.06.2007 20:00(UTC) - 18.06.2007 22:00) Due to minor update to v. 10.1.0.5.0  FTS Database will be affected
CERN 15 JUNE 2007

11:09

Scheduled downtime for CASTORLHCB - Monday June 18 from 07:00 UTC to 16:00 UTC CASTORLHCB instance at CERN will be upgraded to the latest CASTOR release
GRIDKA 14 JUNE 2007

19:18

End of unscheduled maintenance of LFC Problem with the LFC database is fixed
IN2P3 14 JUNE 2007

18:05

IN2P3-CC site will be unavailable from June 20 8:00 to June 21 12:00.   Due to a maintenance of the over all power feed of the building. 
BNL 14 JUNE 2007

17:15

Service intervention at BNL (15.06.2007 12:00 - 15.06.2007 14:00 ) 3D BNL ATLAS service  will be affected
RAL 14 JUNE 2007

13:50

RAL-LCG2 Castor back in service  
GRIDKA 14 JUNE 2007

13:28

Unscheduled maintenance of GridKa LFC  
CERN 14 JUNE 2007

11:29

CERN FTS services will be unavailable for VOs ALICE, ATLAS, CMS, LHCB, DTEAM and OPS. on18.06.2007 from 09:00 to 12:30 prod-fts-ws.cern.ch  & tiertwo-fts-ws.cern.ch to be upgraded to FTS v2.0
RAL 14 JUNE 2007

10:27

RAL-LCG2 CASTOR seems to have reached a state of near meltdown overnight  
CERN 13 JUNE 2007

 09:00

lxplus access problem Fixed at 09:20 hrs.
CERN 13 JUNE 2007

 

The ATLAS CASTOR services at CERN will be unavailable from
07.00 to 16.0 UTC (09.00 to 18.00 CET) on Wednesday 13 June.
Due to upgrade
         
NDGF 12 JUNE 2007

10:00

Service intervention at NDGF from 10:00 on12.06.2007 Affected database: ATLAS (db1tier1.ndgf.org) 
CERN 12 JUNE

09:31

The public CASTOR2 service at CERN will be unavailable from 07.00 to 16.0 UTC (09.00 to 18.00 CET) on Tuesday 12 June TODAY   Due to upgrade
SARA 12 JUNE   Service intervention at SARA saradb will be affected
IN2P3 11 JUNE 2007

18:35

IN2P3-CC available - shutdown postponed Due to a technical problem the shutdown scheduled for tomorrow is postponed.
IN2P3 11 JUNE 2007

17:00

IN2P3-LAPP power cut Unscheduled power cut
IN2P3 11 JUNE 2007

10:00

Service intervention at IN2P3  (13.06.2007 07:00 - 13.06.2007 10:30 ) Affected database: atldb
GRIDKA 10 JUNE 2007   Service intervention at GRIDKA (13.06.2007 07:00 - 13.06.2007 10:30 ) GridKa 3D ATLAS Database will be affected
ASGC 09 JUNE 2007     Service intervention at ASGC (13.06.2007 07:00 - 13.06.2007 10:30 ) ASGC 3D Database will be affected
CNAF-INFN 08 JUNE 2007  

Unexpected down at the INFN Tier1 extended

For the CASTOR service 
RAL 08 JUNE 2007    Service intervention at RAL (13.06.2007 07:00 - 13.06.2007 11:00 )  Database ogma and lugh will be affected
PIC 08 JUNE 2007

17:24

Service intervention at PIC (13.06.2007 07:00 - 13.06.2007 10:30) Service affected  will be the 3D PIC ATLAS service. 
BNL 08 JUNE 2007

17:18

Service intervention at BNL (13.06.2007 07:00 - 13.06.2007 10:30) Service affected  will be the 3D BNL ATLAS service.
IN2P3 08 JUNE 2007

13:47 

CIC Portal is operational  
IN2P3 08 JUNE 2007

12:44

CIC Portal is not reachable. Due to CNAF unschedule downtime
NDGF 08 JUNE 2007

10:06

Service intervention at NDGF (12.06.2007 10:00 - 13.06.2007 14:00 ) Affected database: ATLAS (db1tier1.ndgf.org)
INFN-CNAF 08 JUNE 2007

10:03

INFN-CNAF POWER CUT Power cut since 07.06.2007 night
SARA 08 JUNE 2007

07:43

Service intervention at SARA from 20:40 hrs 07.06.2007  to 08:00 hrs 08.06.2007  
CERN 07 JUNE 2007

14:04

CASTOR2 service at CERN will be unavailable from 07.00 to 16.0 UTC (09.00 to 18.00 CET) on Tuesday 12 June Due to upgrade
CERN 07 JUNE 2007

 

The ATLAS CASTOR services at CERN will be unavailable from
07.00 to 16.0 UTC (09.00 to 18.00 CET) on Wednesday 13 June.
Due to upgrade
RAL 07 JUNE 2007

17:12

Service intervention at RAL on 08.06.2007 12:30 - 08.06.2007 14:30  
SARA 07 JUNE 2007

18:55

On monday june 11th the sara-r1 router will be in maintenance for hardware and software upgrades.  
IN2P3 06 JUNE 2007

08:44

MSS (HPSS) is back online  
IN2P3 06 JUNE 2007

11:32

[IN2P3-CC] SRM unscheduled downtime Due to unexpected problem on  SRM server
SARA 06 JUNE 2007

12:52

Maintenance SARA LFC_oracle and FTS  
IN2P3 06 JUNE 2007

15:36

SRM server is back online    
RAL 06 JUNE 2007

16:16

Short downtime on RAL-LCG2 CE on 7 June 2007 2pm GMT  
SARA 06 JUNE 2007

16:40

LFC_oracle and FTS at SARA is back online  
IN2P3 06 JUNE 2007

19:07

Network problem between IN2P3-CC and CERN Due to a router configuration error at IN2P3-CC
IN2P3 06 JUNE 2007

21:43

Problem between IN2P3-CC and CERN SOLVED  
IN2P3 05 JUNE 2007

13:16

CIC portal back online from INFN-CNAF  
RAL 05 JUNE 2007

13:17

RAL lcg-RB scheduled maintenance finished  
RAL 05 JUNE 2007

14:11

RAL-LCG2 dcache-tape.gridpp.rl.ac.uk tape backend problem Due to a problem with the configuration of the Fibre Channel switches
RAL 05 JUNE 2007

16:02

RAL-LCG2 dcache-tape backend available  
RAL 05 JUNE 2007

16:31

Service intervention at RAL on 07.06.2006  
IN2P3 05 JUNE 2007

18:28

HPSS problem at IN2P3-CC.  
SARA 04 JUNE 2007

07:46

Service intervention at SARA A module in one of the SARA routers crashed
NDGF-T1 04 JUNE 2007

08:53

NDGF-T1 FTS inactive at the moment Due to database space problems
CERN 04 JUNE 2007

09:38

All Atlas data available at IN2P3-CC SRM  
NDGF-T1 04 JUNE 2007

12:22

NDGF-T1 FTS back online  
CERN 04 JUNE 2007

14:28

BDII failure on ce122.cern.ch -(CERN-PROD)  
FZK / GridKA 04 JUNE 2007

15:52

Network Problems at GridKa  
IN2P3 04 JUNE 2007

16:36

IN2P3-CC unavailable from june 11th at 8:00 to june 13th at 12:00 Due to scheduled electrical shutdown
FZK / GridKA 04 JUNE 2007

16:48

Network Problems at GridKa Fixed  
IN2P3 02 JUNE 2007   Part of Atlas data unavailable at IN2P3-CC SRM Due to a hardware problem on one of the SRM-dCache pool nodes
CERN 01 JUNE 2007

10:43

Problem with SAM tests causing all sites to fail Due to problem with permissions on LFC at CERN
RAL 01 JUNE 2007

14:10

RAL Tier 1 database OGMA under maintenance Unscheduled maintenance
SARA 01 JUNE 2007

16:54

Network problem at SARA  
FZK / GridKA 01 JUNE 2007   Network interrupts on 8 June 2007 Due to maintenance on the routers
CERN 31 MAY 2007

08:42

CE failure on ce101.cern.ch -(CERN-PROD)-  
CNRS 31 MAY 2007

10:56

CNRS GRID-FR CA will be down on Saturday, June 9th 2007, from 7h30 am CEST to 11h30 am CEST. For scheduled electric power maintenance
IN2P3  31 MAY 2007

17:26

CIC portal will be unavailable on Tuesday 5th from 07:00 UTC to 12:00 UTC (09:00 CET to 12:00 CET),  
CERN 31 MAY 2007   LCG RB rb107.cern.ch dedicated to CMS down  
RAL 31 MAY 2007

17:32

Castor unavailable  
SARA 31 MAY 2007

18:57

SARA LFC_mysql is back again  
SARA 30 MAY 2007

08:55

Scheduled maintenance for oracle server at SARA  
CERN 30 MAY 2007

10:17

Problems with CERN-PROD/castorcms last night Due to garbage collection
CERN 30 MAY 2007

11:20

Sms down Due to CERNDB1 down
NDGF-T1 30 MAY 2007

16:08

NDGF-T1 SRM service stop  extended  
NDGF-T1 30 MAY 2007

18:25

NDGF-T1 SRM service available  
CERN 30 MAY 2007

 

CERN LCG CEs problem

Due to high activity
CERN 29 MAY 2007

09:04

SAM Admin''s Page  not responding Due to SAM client problems with connections to SAM framework
CERN 29 MAY 2007

14:54

UPDATE 25 for gLite 3.0  released to the EGEE Grid production service.  
PragueLCG2 25 MAY 2007

09:27

Praguelcg2 site down till Monday praguelcg2 power failure
CERN 25 MAY 2007

14:05

SAM service back in production  
NDGF-T1 25 MAY 2007

16:03

 NDGF-T1 SRM service will stop  from 08:00 UTC (10:00 CET) to approx. 12:00 UTC.on Wednesday, May 30th Due to a software upgrade

 

NDGF-T1 25 MAY 2007

16:42

 NDGF-T1 FTS service will be running "at risk" on Sunday. Due to a planned 8h power outage
RAL-LCG2 24 MAY 2007

10:22

RAL-LCG2 Castor currently unavailable  
SARA 24 MAY 2007

10:36

LFC problems at SARA  
CERN 24 MAY 2007

10:42

lcg-voms.cern.ch certificate has been changed  
RAL-LCG2 24 MAY 2007

12:06

RAL-LCG2 Castor now available  
RAL-LCG2 24 MAY 2007

12:53

Possible disruptions with R-GMA registry at RAL-LCG2  
SARA 24 MAY 2007

15:29

LFC at SARA is back in operation  
CERN 24 MAY 2007

16.35

New release of the GGUS portal online  
CERN 24 MAY 2007

18:01

Availability problems on VOMS servers Due to problems on the OracleDB backend
CERN 24 MAY 2007

21:42

Unscheduled downtime of SAM at CERN Because of a load of unknown origin caused on the SAM DB
RAL-LCG2 23 MAY 2007

10:38

The RAL-LCG2 FTS will be unavailable between 09:00 and 13:00 GMT+1 on May 30 Oracle and kernel updates
SARA 23 MAY 2007

15:26

LFC at SARA is back online  
CERN 23 MAY 2007

15:32

A new version of the UI has been installed on the CERN AFS space, version 3.0.21-0 corresponding to UPDATE 24.  
CERN 22 MAY 2007

11:04

 LAST WARNING: lcg-voms.cern.ch certificate will be changed on May 24th!  
PragueLCG2 22 MAY 2007

12:21

Data lost on se2.farm.particle.cz The raid controller and disks have failed on one of  DPM pool nodes
CERN 22 MAY 2007

12:43

Glite-WN software is being upgraded on LXBATCH/SLC3 at CERN  
RAL-LCG2 22 MAY 2007

15:18

A maintenance operation is scheduled for the second Resource Broker at RAL-LCG2 (lcgrb02.gridpp.rl.ac.uk) - starting with Tuesday 29 May 08:00 UTC (9am BST) The machine will be hopefully back on line and accepting job submissions on Wednesday 6 June 15:00 UTC (4pm BST)
CERN 22 MAY 2007

15:31

gLite restructuring plan to kick off on May 30 2007  
IN2P3 22 MAY 2007

17:36

IN2P3-CC unavailable on 28-29/05/2007 Hardware upgrade for  local batch system
CERN 22 MAY 2007

19:50

All gLite WMS nodes at CERN back in production  
IN2P3 21 MAY 2007   LCG RB grid09.lal.in2p3.fr has been shutdown today  
FZK/GridKa 21 MAY 2007

19:38

ce-gridka.fzk.de rebooted  
SARA 18 MAY 2007

09:16

Disk problem of Oracle server at SARA  
UAM-LCG2 18 MAY 2007

10:17

Scheduled Downtime (UAM-LCG2) Start: 21st May 2007 - 08:00 End: 25th May 2007 - 14:00  Maintenance of Hep Laboratory Electrical Power (add UPS)
Belgrid-UCL 18 MAY 2007

11:11

Scheduled Dowtime at Belgrid-UCL  Begin: 21st May 2007 - 08:00 End: 25th May 2007 - 20:00  Network configuration update, PostgreSQL server upgrade, WN update
TRIUMP 17 MAY 2007

05:52

TRIUMF dCache pool dead  
TRIUMP 17 MAY 2007

07:34

dCache pool back in service  
SARA 17MAY 2007

10:38

 Internal networking problems at SARA  
RAL 17 MAY 2007

14:31

RAL-LCG2 Castor now available  
SARA 16 MAY 2007

09:46

Problems with the LFC on mu11.matr.sara.nl and the FTS. Problems with the oracle server
CERN 16 MAY 2007

10:31

GGUS system will be down from 12:00 to 16:00 UTC on 24th. May The new release of the GGUS portal will be made available on Thursday, 24th of May 2007.
SARA 16MAY 2007

11:15

Unscheduled maintenance at SARA Due to internal networking problems SARA-MATRIC and SARA-LISA
CERN 16MAY 2007

11:23

gCE failure on ce109.cern.ch  
CERN 16 MAY 2007

13:05

Bug found in LFC/DPM 1.6.4-3, distributed in gLite Update 24  
SARA 16 MAY 2007

19:26

SARA internal networking problems  
CERN 15 MAY 2007  13:09 LCG - VOMS.Cern.ch certificate change on MAY 24 2007 Warning
CERN 15 MAY 2007  13:43 New Version (3.0.14-4) of CERN AFS UI for CERN PROD  
CERN 15 MAY 2007 14.18 Storage01-lcg.projects.cscs.ch down Problem with dcache
CERN 15 MAY 2007  15:40 UPDATE 24 for gLite 3.0  released to the EGEE Grid production service.